# Nova: Generative Language Models for Assembly Code with Hierarchical Attention and Contrastive Learning

### v4

@@ 6 Conclusion
7 Limitations
One limitation is that Nova is X86-specific, as we only collect X86 assembly corpus for pre-training. This design choice is mainly affected by two facts: (1) X86 assembly is used and explored in a wide range of binary tasks Wang et al. (2022); Su et al. (2024); Xu et al. (2023a); Chen et al. (2022c) compared to other assembly languages, and (2) computation limitations. However, the proposed techniques are independent of X86 assembly. Low information density and compiler optimization are the common challenges of most assembly languages such as X86, ARM, and MIPS. The proposed techniques can be applied to the future development of multi-lingual assembly LLMs.
@@ A.1 Data Collection
A.2 Hierarchical Self-Attention
The hierarchical self-attention is designed for assembly code, yet the input to LLMs may still contain text or source code. Figure 8 illustrates how the hierarchical attention works with text or source code in the input. As existing LLMs have shown good performance on text and source code using the standard self-attention, we keep the standard causal attention mask within and between any chunks of text or source code in the input (the light grey part shown in Figure 8).
The attention from text or source code to assembly code (and vice versa) is only allowed for the “[INST]” tokens as they are supposed to contain the assembly instruction summaries.
Figure 8: Hierarchical attention with text input.
@@ A.2 Hierarchical Self-Attention
A.3 Training Details
The fine-tuning of both BCD and BCSD uses a batch size of 64, with the input truncated by a 2,048 token limit. Similarly, the learning rate is $2e^{-5}$ using the AdamW optimizer, and the model is fine-tuned for one epoch.
Figure 9: Nova-1B correctly decompiles HumanEval-Decompile task 0, while GPT-4 and LLM4Decompile-1B fail.
Figure 10: Nova-1B correctly decompiles HumanEval-Decompile task 20, while LLM4Decompile-1B fail.
Figure 12: t-SNE of embeddings cauclated by Nova-CL-HA, Nova-HA, and Nova.
@@ A.5 Binary Code Similarity Detection Ablation Study
Table $13$ show the detailed ablation study results of BCSD. Nova wins on the most benchmarks when $K=100$ or $500$, and ties with Nova-HA when $K=50$, or $200$.
@@ A.6 Additional Analysis of Embedding
Additional Embedding Visualization
Figure 11 shows the full results of PCA of embeddings provided by Nova-CL-HA, Nova-HA, and Nova, on randomly sampled seven examples. Compared with Nova-CL-HA, Nova-HA including contrastive learning objectives in the pre-training, can separate the embeddings of assemblies with different functionalities better. Nova-HA clearly encode “Task 143” (orange points) away from the others. Nova’s embeddings group the assemblies by functionalities more precisely than Nova-HA, suggesting that hierarchical attention enhances the training of contrastive learning objectives to learn more effective encoding.
Figure 12 shows the results using another dimensionality reduction technique, t-SNE van der Maaten and Hinton (2008), where Nova’s embeddings are consistently more distinguishable by functionalities.
Figure 13: Distances between embeddings of different-formats-same-functionality code and same-format-different-functionalities code.
Figure 14: Comparison of attention distribution among standard and hierarchical heads in the final layer.
Quantitative Analysis
We also perform quantitative analysis to further support the conclusion from visualization. Figure 13 (a) shows the distribution of $l_{2}$ distances between different-formats-same-functionality code in HumanEval-Decompile (“format” is defined as one of source code, O0, O1, O2, or O3 assembly). The figure shows five distributions, (src, O0), (src, O1), (src, O2), and (src, O3). Each distribution calculates the distances between embeddings of two formats of code, e.g., src, O0 refers to the source code and O0 assembly for the same task in HumanEval-Decompile.
Figure 13 (a) shows that Nova’s embeddings for same-functionality-different-formats code are closer to each other in the latent space. For the same task, all the O0O3 assemblies’ embeddings have a smaller $l_{2}$ distance to the source code embedding compared to that of Nova-CL-HA’s and Nova-HA’s embeddings. Another interesting finding is that for Nova, the distances between assembly and source code increase with the optimization level of assemblies increases. This trend matches what the optimization CL tries to optimize for.
Figure 13 (b) shows the distribution of $l_{2}$ distance between same-format-different-functionalities code from the HumanEval-Decompile benchmark. With the same format (source code or the same optimized assembly), Nova encodes the assemblies with different functionalities farther away from each other compared to Nova-CL-HA. Nova-CL-HA’s embeddings cannot significantly reflect the functionality differences between assemblies, while Nova’s embeddings can.
Nova-HA’s embeddings show high $l_{2}$ distance in both Figure 13 (a) and (b), suggesting that Nova-HA tries to decentralize all the code in the embedding space even if they have the same functionality, which is not as desired as Nova.
@@ A.7 Additional Analysis of Attention
In addition, in lower layers, we have observed attention weights to be softly distributed among tokens comprising each instruction (Figure 15), which suggests Nova initially models cross-relations among operation codes and operands in the first few layers, and later pools their summary representation into the [INST] token in the later layers. This is also supported by Figure 6, where the Nova’s hierarchical attention (red line) shows a decreasing trend of entropy. This means the hierarchical attention is softer in lower layers, and becomes focused in higher layers.

### v5

@@ ACKNOWLEDGMENTS
This paper was prepared for informational purposes in part by the CDAO group of JPMorgan Chase & Co and its affiliates (”J.P. Morgan”) and is not a product of the Research Department of J.P. Morgan. J.P. Morgan makes no representation and warranty whatsoever and disclaims all liability, for the completeness, accuracy or reliability of the information contained herein. This document is not intended as investment research or investment advice, or a recommendation, offer or solicitation for the purchase or sale of any security, financial instrument, financial product or service, or to be used in any way for evaluating the merits of participating in any transaction, and shall not constitute a solicitation under any jurisdiction or to any person, if such solicitation under such jurisdiction or to such person would be unlawful.
@@ A.1 Data Collection
A.2 Training Details
The fine-tuning of both BCD and BCSD uses a batch size of 64, with the input truncated by a 2,048 token limit. Similarly, the learning rate is $2e^{-5}$ using the AdamW optimizer, and the model is fine-tuned for one epoch. During the training using the contrastive learning objectives, and the fine-tuning of BCSD, we use the average of [INST] tokens’ last layer hidden states to represent the embedding of a binary function.
Table 10: Comparison with applying hierarchical attention on all attention heads, using 1B models.
Techniques Pass@1 Pass@10
O0 O1 O2 O3 Avg. O0 O1 O2 O3 Avg.
Nova-HA 30.58 19.88 20.58 16.40 21.86 44.75 33.13 33.31 29.82 35.25
Nova (hierarchical on all heads) 32.38 18.87 20.56 16.34 22.04 45.95 32.19 32.78 29.01 34.98
Nova 37.53 21.71 22.68 18.75 25.17 49.38 34.84 36.95 32.03 38.30
A.3 Applying Hierarchical Attention on Half Attention Heads
The hierarchical attention mask is always applied on half of the attention heads at each layer in Nova. This ensures the LLM balances the hierarchical knowledge of assembly code and pre-trained knowledge learned by full self-attention.
Figure 7: Nova-1B correctly decompiles HumanEval-Decompile task 0.
We conducted experiments applying hierarchical attention to all the attention heads. Results in Table 10 show that when applying hierarchical attention to all the attention heads of transformer layers, the performance does not improve enough and even drops under some settings. This implies that the standard full self-attention mechanism indeed learns knowledge that may not be captured by hierarchical attention. Thus, we only apply hierarchical attention to half of the attention heads in each transformer layer of Nova to balance the knowledge learned by standard and hierarchical attention.
Figure 8: Nova-1B correctly decompiles HumanEval-Decompile task 20.
@@ A.4 Binary Code Decompilation Case Studies
Figure 10: Comparison of attention distribution among standard and hierarchical heads.
@@ A.5 Binary Code Similarity Detection Ablation Study
Table $14$ show the detailed ablation study results of BCSD. Nova wins on the most benchmarks when $K=100$ or $500$, and ties with Nova-HA when $K=50$, or $200$.
@@ A.6 Additional Analysis of Embedding
Figure 9 shows the results of PCA of embeddings provided by Nova-CL-HA, Nova-HA, and Nova, on randomly sampled seven examples, where Nova’s embeddings are consistently more distinguishable by functionalities.
@@ A.7 Additional Analysis of Attention
In addition, in lower layers, we have observed attention weights to be softly distributed among tokens comprising each instruction (Figure 11), which suggests Nova initially models cross-relations among operation codes and operands in the first few layers, and later pools their summary representation into the [INST] token in the later layers.